OpenStudy (anonymous):

The side of a pyramid is 50m long and forms a 30 degree angle with the ground. What is the height of the pyramid?

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

sine uses opposite and hypotenuse sin(30) = opp/hyp sin(30) = h/50 1/2 = h/50 ... use a calculator to get sin(30) = 1/2 h = ??

OpenStudy (anonymous):

I don't understand...

jimthompson5910 (jim_thompson5910):

where are you stuck?

OpenStudy (anonymous):

when i do sin(30) its negative....

jimthompson5910 (jim_thompson5910):

you need to be in degree mode (not radian mode)

OpenStudy (anonymous):

oh duh! thanks!

jimthompson5910 (jim_thompson5910):

you're welcome